OpenStudy (anonymous):

In a paragraph, describe how to simplify each expression. 16 1/2 √100

OpenStudy (anonymous):

16 1/2 = 16+1/2=[16(2)+1]/2= 33/2 sqrt100=sqrt(10x10)=10

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

YWC